8) 23.87: An electric dipole in a uniform horizontal electric field is displaced slightly from its equilibrium position as shown, where e is small. The separation of the charges is 2a, and each of the two particles has mass m. a)Assuming the dipole is released from this position (at res), show that its angular orientation exhibits simple harmonic motion (this is the same as showing that d +00 0-0 where a is a collection of constants; this may be done by writing the torque d t equivalent of Newtons 2nd Law and rearranging) with a frequency f b) Suppose the 2 ma
Answer #1

Torque exerted by electrostatic force on the positive charge

\large \tau = \vec{r} \times \vec{F}

\large \tau = -rF\sin \theta

Here electrostatic force F = qE, r = a Taking centre of the dipole as axis of rotation

\large \tau = -aqE\sin \theta

Similarly torque on negative charge

\large \tau = -aqE\sin \theta

Net Torque

\large \tau _{net}= -2aqE\sin \theta

From Newton's Second law of rotational Motion

\large \tau_{net} = I\alpha = I\frac{\mathrm{d^{2}} \theta }{\mathrm{d} t^{2}}

Here Moment of inertia of the dipole is given by

\large I = ma^{2}+ma^{2} = 2ma^{2}

Using all the equations we get

\large 2ma^{2}\frac{\mathrm{d^{2}} \theta }{\mathrm{d} t^{2}} = -2aqE\sin \theta

\large ma\frac{\mathrm{d^{2}} \theta }{\mathrm{d} t^{2}} = -qE\sin \theta

\large \frac{\mathrm{d^{2}} \theta }{\mathrm{d} t^{2}} = -\frac{qE}{ma}\sin \theta

For small angular displacement \large \sin \theta \approx \theta

\large \frac{\mathrm{d^{2}} \theta }{\mathrm{d} t^{2}} = -\frac{qE}{ma} \theta

So we can say that

\large \frac{\mathrm{d^{2}} \theta }{\mathrm{d} t^{2}} \propto \theta

So Dipole will execute the SHM.

Now comparing with standred equation of SHM we get

\large \omega ^{2} = \frac{qE}{ma}

\large \omega = \sqrt{\frac{qE}{ma} }

We know that

\large \omega = 2\pi f= \sqrt{\frac{qE}{ma} }  So

\large f= \frac{1}{2\pi }\sqrt{\frac{qE}{ma} }